A child is pushing a merry-go-round. The angle through which the merry-go-round has turned varies with time according to θ(t)= γt+ βt3, where γ = 0.406 rad/s and β = 1.30×10−2 rad/s3.

a) Calculate the angular velocity of the merry-go-round as a function of time.

b) What is the initial value of the angular velocity?

c) Calculate the instantaneous value of the angular velocity ωz at t=0 to t= 5.00 s. Show that the wav-z is not equal to the average of the instantaneous angular velocities at t=0 and t=5.00 s, and explain why it is not.


1
Expert's answer
2022-01-20T10:07:43-0500

"\text{ a)}\space \omega(t)=\theta'(t)=\gamma+3\beta t^2.\\
\text{ b)}\space \omega(0)=\gamma=0.406 \text{ rad/s}.\\
\text{ c)}\space \omega_\text z(5)=0.406 +3·0.012·5^2=1.38\text{ rad/s}.\\
\space\space\space\space\space\space\omega_\text{av}=[\omega_\text z(5)-\omega(0)]/t,\\
\space \space\space\space\space\space\omega_\text{av}=[1.38-0.406]/5=0.195\text{ rad/s}."

The average speed is different because we use a different formula to calculate it.